Given the below info, about an OPEN economy, calculate the equilibrium level of income?

Consumption function =100+0.7YD. Investment = 200. Govt spending =15. Tax = 0.2 Exports 100. Imports 100

Answer is 859
But can you help explain how they get to 859
1
Expert's answer
2019-04-08T09:47:18-0400

Solution

YD=Y-T

Equilibrium level of income=C+I+G+(E-I)=100+0.7(Y-0.2Y)+200+15+(100-100)=Y

315+0.56Y=Y

315=0.44Y

Y=715.91

100+0.7*(71.91-0.2*71.91)+200+15+0=715.91

Answer is 715.91.